A 10,000 kg railroad car is rolling at 2.00 m/s when a 2000 kg load of gravel is suddenly dropped in. What is the car's speed just after the gravel is loaded?
Answer #1

Solution)

Mass, m1=10,000 kg

Velocity, v1=2 m/s

Mass, m2=10,000 + 2000=12,000 kg

Velocity, v2=?

According to Conservation of Linear Momentum,

M1v1= m2v2

Substitute values,

10,000*2=12000*v2

So, speed, v2= 1.66 m/s (Ans)

===========